OpenStudy (anonymous):

-3 over 6 + 7 over 12 =?

OpenStudy (anonymous):

\[-\frac{ 3 }{ 6 } +\frac{ 7 }{ 12 } =?\]

OpenStudy (anonymous):



OpenStudy (anonymous):

that doesnt make sense to me.

OpenStudy (anonymous):

1/12

OpenStudy (anonymous):

-36/13

OpenStudy (smored):

you need the LCD (least common denominator) of 6 and 12, which is 12. \[\frac{ -3 }{ 6 }+ \frac{7}{12}\] equals \[\frac{ -6 }{ 12 }+ \frac{7}{12}\] which equals 1/12

OpenStudy (smored):

@chetan552 that makes no sense.